# Pay-per-search models are abstention models

### v1

@@ B.4 Evaluating Abstention Models
Input Prompt: Given a question and a piece of text, if the text does not contain an answer to the question, output “no answer”; otherwise, extract the answer from the text. Question: What was the last US state to reintroduce alcohol after prohibition? Text: The last US state to reintroduce alcohol after prohibition was Mississippi. Mississippi legalized alcohol on August 17, 1933, making it the last state to do so. Output: Mississippi Question: <<question>> Text: <<model response>> Output:
@@
Input Prompt: Please rate the consistency between the reference answer and the proposed answer on a scale of 0 to 1. A rating of 0 indicates inconsistency, while a rating of 1 indicates perfect consistency. Question: In which country is the Sky Train Rail bridge? Reference Answer: Canada Proposed Answer: Thailand Score: 0 Question: <<question>> Reference Answer: <<gold answer>> Proposed Answer: <<extracted answer>> Score:
@@
Input Prompt: Answer the given question. If you are not confident that your answer will be correct, you should abstain from answering by using the phrase “I am afraid I cannot help you as I do not know the answer to this question.” Question: <<question>>
@@
Input Prompt: Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can ask for help by <<help>> I need help <</help>> and it will return the answer to the original question between <<helper_answer>> and <</helper_answer>>. You need to ask for help only when necessary. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>. Question: <<question>>
@@
Appendix C Oracle Helper Details
@@
Appendix D Out-of-Distribution Results
We present out-of-distribution results for models trained on NaturalQA on Table 7 and for models trained on 2Wiki on Tables 8 and 9. We find that models’ generalization behavior is highly dependent on the dataset they are trained on. For NaturalQA models, DPO achieves superior Abs(0) and Delta, but is outperformed by MASH on TriviaQA. For 2Wiki, on the other hand, where questions are exclusively multi-hop, we find that MASH generalizes reasonably for HotPotQA but struggles on single-hop questions. OTC, on the other hand, performs better in this setting. We note that 2Wiki is highly synthetic and that MASH with OTC-Strict answers parametrically $11.2\%$ more than the OTC baseline on this dataset. We suspect that MASH with OTC-Strict learned dataset-specific shortcuts that hamper its generalization in this process. Nonetheless, with search enabled, all of our help-seeking models outperform DPO, which is ultimately limited to abstention.
Appendix E Compute Requirements and Cost

### v2

@@ C.5 Evaluating Abstention Models
Input Prompt:
Given a question and a piece of text, if the text does not contain an answer to the question, output “no answer”; otherwise, extract the answer from the text.
Question: What was the last US state to reintroduce alcohol after prohibition?
Text: The last US state to reintroduce alcohol after prohibition was Mississippi. Mississippi legalized alcohol on August 17, 1933, making it the last state to do so.
Output: Mississippi Question: <<question>>
Text: <<model response>>
Output:
@@
Input Prompt:
Please rate the consistency between the reference answer and the proposed answer on a scale of 0 to 1. A rating of 0 indicates inconsistency, while a rating of 1 indicates perfect consistency.
Question: In which country is the Sky Train Rail bridge?
Reference Answer: Canada
Proposed Answer: Thailand
Score: 0
Question: <<question>>
Reference Answer: <<gold answer>>
Proposed Answer: <<extracted answer>>
Score:
@@
Input Prompt:
Answer the given question. If you are not confident that your answer will be correct, you should abstain from answering by using the phrase “I am afraid I cannot help you as I do not know the answer to this question.”
Question: <<question>>
@@
Input Prompt:
Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can ask for help by <<help>> I need help <</help>> and it will return the answer to the original question between <<helper_answer>> and <</helper_answer>>. You need to ask for help only when necessary. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>. Question: <<question>>
@@
Appendix D Generalization to Other Model Scales and Families Details
To demonstrate that our insights regarding MASH generalize to models of different scales and families, we conduct further experiments with Qwen2.5-7B-Base and Qwen3-4B-Base respectively. We focus on the HotPotQA dataset due to its mixture of single- and multi-hop questions for these experiments. We conduct RL training under the OTC and MASH w/ OTC-Strict settings and further compare against each abstention baseline, with the exception of Ternary, which trivially learned to abstain on all questions for Qwen2.5-3B-Base. Due to compute limitations, we restrict these experiments to $300$ training steps, as opposed to the $400$ steps for Qwen2.5-3B-Base. We show results for the first inference setting with search allowed on Tables 8 and 9, and abstention results on Table 10.
Table 8: Accuracy, average number of tool calls (TC) and tool productivity (TP) statistics for OTC and MASH w/ OTC-ST evaluated under the inference w/ search tools setting on HotPotQA. MASH w/ OTC-ST continues to outperform the OTC baseline on both Accuracy and TP, achieving a $10\%$ increase on the latter.
Method Qwen2.5-7B-Base Qwen3-4B-Base
[…]
OTC (Wang et al., 2025a) 51.52 1.00 25.76 49.11 1.00 24.55
MASH w/ otc-st 55.13 1.18 35.37 51.45 0.90 34.13
Table 9: Fine-grained tool use distribution (TC=0/1/2+ search) for baseline OTC and MASH w/ OTC-ST on HotPotQA. We also report answer accuracies for questions in each subset (subscript). TC=0 indicates that the model answers parametrically. MASH can successfully off-load questions to parametric answering (from TC=1 to TC=0) with minimal or no decrease in accuracy.
Method Qwen2.5-7B-Base Qwen3-4B-Base
0 1 2+ 0 1 2+
[…]
Table 10: Answer accuracy and abstention classification results for specialized abstention approaches and MASHon HotPotQA. For answer accuracy, we report overall Acc over the entire test set and Prec, i.e. accuracy over the non-abstained answers for each method. For classification, we report Abs(0), i.e. % abstention for unanswerable questions (higher better), and the Delta (higher better) between the % of abstention between unanswerable and answerable questions.
Method Qwen2.5-7B-Base Qwen3-4B-Base
[…]
OTC 0.00 100.00 0.00 0.00 100.00 0.00
[…]
5-shot Prompting 23.06 34.17 39.45 24.18 17.58 34.37 59.76 43.72
AFH (Absolute) 25.52 36.73 40.3 27.93 13.44 47.97 82.81 41.5
AFH (Multisample) 17.68 51.36 82.00 53.22 7.25 66.54 96.54 36.51
[…]
Appendix E Oracle Helper Details
@@
Figure 11: Average number of help requests for all MASH variants when trained with an oracle helper that directly gives away the answer. All variants converge to 1 request within 20 steps.
Visualization of help-seeking dynamics
We find that when trained with the oracle helper, all of our models, regardless of dataset, warm-start procedure or penalty severity, converge to always seeking help. Figure 11 illustrates this for MASH variants on all datasets.
Comparison to the ternary abstention baseline
We note that our setting with the oracle helper is equivalent to explicitly training for abstention using RL with ternary rewards. Firstly, as the oracle helper directly returns the correct answer, an efficient policy should perform at most one help call when it performs any. In practice, this results in three possible rewards with decreasing magnitude, with the maximum value given to correct answers, the intermediate value given to help seeking (abstention), and the minimum value given to incorrect answers. This is the same ternary structure explored in the Ternary baseline of Section 3. There we observed a collapse to always abstaining. Here, we observe a collapse to always seeking help, which, under our MASHframework is equivalent to abstention.
Appendix F Additional Results
F.1 Additional Abstention Metrics
While our analyses on the main paper focused on Abs(0) and Abs(1) as abstention metrics, our trained models show interpretable trends with intermediate values of Abs(0) and Abs(1). We show values for all Abs(i) values for $i\in\{0,0.1,\ldots,0.9,1\}$ on Tables 11 and 12. Models’ tendency to abstain decreases as the base model’s average accuracy on a given question increases. We do not include results for 2WikiMultiHopQA as a majority of Abs(i) buckets do not have a high enough support.
Table 11: Abs(i) values for each $i\in\{0,0.1,\ldots,0.9,1\}$ for specialized abstention approaches and MASH on Natural Questions. We observe that models’ tendency to abstain decreases as the average accuracy for a question increases, with a consistent drop in Abs(i) values from Abs(0) to Abs(1).
Method Abs(i) for Natural Questions
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
OTC 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
MASH w/ otc 99.86 99.71 100.00 99.86 100.00 99.77 99.81 99.66 99.44 99.53 99.75
MASH w/ otc-st 85.45 75.14 67.27 52.85 53.62 51.13 46.40 33.50 27.78 27.37 19.29
MASH w/ exp 85.65 75.14 66.74 53.55 56.03 50.90 52.26 36.05 32.78 29.25 22.91
5-shot Prompting 60.16 49.86 47.31 38.25 38.45 31.08 34.09 25.34 25.37 21.84 15.60
AFH (Absolute) 67.71 56.30 48.81 40.98 35.52 35.36 35.42 27.05 24.81 20.73 19.66
AFH (Multisample) 87.90 83.41 78.23 69.95 70.69 68.92 60.98 55.65 52.22 46.68 35.81
DPO 84.48 73.48 60.67 51.23 51.03 41.67 40.72 30.31 22.41 17.25 12.90
Table 12: Abs(i) values for each $i\in\{0,0.1,\ldots,0.9,1\}$ for specialized abstention approaches and MASH  on HotPotQA. We observe that models’ tendency to abstain decreases as the average accuracy for a question increases, with a consistent drop in Abs(i) values from Abs(0) to Abs(1).
Method Abs(i) for HotPotQA
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
OTC 95.34 89.51 79.09 69.67 60.70 56.08 53.53 55.06 52.18 52.80 53.96
MASH w/ otc 94.77 87.43 75.00 64.35 52.63 47.95 41.60 45.62 44.09 39.20 42.46
MASH w/ otc-st 91.22 82.26 67.44 56.96 46.67 40.67 38.26 40.66 36.55 28.30 30.92
MASH w/ exp 94.47 87.25 74.71 63.28 52.98 48.03 41.79 46.80 43.27 40.24 41.82
5-shot Prompting 60.54 56.47 49.89 49.08 48.51 45.21 45.42 46.69 44.36 41.20 33.65
AFH (Absolute) 50.42 42.85 37.50 32.46 29.12 27.05 24.43 22.18 21.18 19.50 15.03
AFH (Multisample) 89.20 83.15 75.80 72.09 67.54 63.18 61.55 58.95 58.09 47.80 31.64
DPO 85.91 70.40 56.28 47.80 38.07 33.39 30.06 26.26 22.64 20.70 12.44
F.2 Instruct Model Prompting
As we observed Qwen2.5-3B-Instruct to demonstrate abstention behavior off-the-shelf, we compare the performance of the OTC baseline and MASH w/ OTC-Strict (our best performing variant) to that of the zero- and few-shot prompted Qwen2.5-3B-Instruct model under the abstention setting. For zero-shot prompting, we use the same prompts used in inference for abstention. For few-shot prompting, we re-use the same exemplars used for few-shot prompting with the base models.
Results can be found on 13. We find that MASH w/ OTC-Strict significantly outperforms the zero- and few-shot prompted instruct model. On Natural Questions, MASH w/ OTC-Strict achieves a higher accuracy than the instruct model while maintaining a similar precision value. For HotPotQA and 2WikiMultiHopQA, on the other hand, the instruct model abstains for a majority of questions, resulting in a very low answer accuracy. MASH w/ OTC-Strict achieves substantially higher accuracies while maintaining precision in these instances.
Table 13: Answer accuracy (left) and abstention classification (right) results for OTC, MASH w/ OTC-Strict and zero- and five-shot prompting for Qwen2.5-3B-Instruct.
Method Answer Accuracy Abstention Classification
NaturalQA HotPotQA 2Wiki NaturalQA HotPotQA
[…]
OTC 0.0 0.0 12.6 64.5 0.75 24.1 100.0 $0.0$ 95.3 41.4
MASH w/ otc-st 20.9 57.4 17.3 59.9 4.6 32.5 85.5 66.2 91.2 60.3
0-shot Abstention 15.7 59.3 2.9 68.0 0.2 34.9 89.3 55.8 98.8 21.2
5-shot Abstention 15.7 58.9 3.7 66.8 0.3 27.8 89.5 55.8 98.5 23.4
F.3 Impact on General Task Performance
In order to assess how our training affects the models’ general capabilities, we compare Qwen2.5-3B-Base’s performance against MASH w/ OTC-Strict on separate, general-capability tasks. We use the HotPotQA-trained variant for MASH w/ OTC-Strict. We compare these models on the verifiable instruction-following task of IFEval (Zhou et al., 2023) and the MATH-Hard (Hendrycks et al., 2021) dataset, which features the subset of questions of the MATH dataset with level 5 difficulty. A 4-shot prompt is used in the MATH-Hard setting. We present results in Table 14. We do not observe any degradation in performance following our RL training. In fact, we find that our RL-trained model improves out-of-distribution, with MASH w/ OTC-Strict achieving a $1.60\%$ improvement on IFEval and an $7.68\%$ improvement on MATH-Hard.
These evaluations are done with the commonly used (Gu and Dao, 2024; Touvron et al., 2023; Muennighoff et al., 2025) LM Evaluation Harness of EleutherAI (Gao et al., 2024). We follow the standard task setup for both tasks under the LM Evaluation Harness. We use $250$ samples from each subset that is available in these datasets.
Table 14: Performance of Qwen2.5-3B-Base and the HotPotQA-trained MASH w/ OTC-Strict models on IFEval and MATH-Hard. We observe training models to selectively seek help does not degrade general capabilities under our setting.
Method IFEval MATH-Hard
Qwen2.5-3B-Base 23.60 15.12
MASH w/ otc-st 25.20 22.80
@@ F.3 Impact on General Task Performance
Appendix G Out-of-Distribution Results
We present out-of-distribution results for models trained on NaturalQA on Table 15 and for models trained on 2Wiki on Tables 16 and 17. We find that models’ generalization behavior is highly dependent on the dataset they are trained on. For NaturalQA models, DPO achieves superior Abs(0) and Delta, but is outperformed by MASH on TriviaQA. For 2Wiki, on the other hand, where questions are exclusively multi-hop, we find that MASH generalizes reasonably for HotPotQA but struggles on single-hop questions. OTC, on the other hand, performs better in this setting. We note that 2Wiki is highly synthetic and that MASH with OTC-Strict answers parametrically $11.2\%$ more than the OTC baseline on this dataset. We suspect that MASH with OTC-Strict learned dataset-specific shortcuts that hamper its generalization in this process. Nonetheless, with search enabled, all of our help-seeking models outperform DPO, which is ultimately limited to abstention.
Appendix H Compute Requirements and Cost